OpenStudy (mikezack123):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of 1/2.

OpenStudy (mikezack123):

y − 6 =1/2 (x + 3) y = 1/2x − 6 y + 3 = 1/2(x − 6) x − 2y = 12

OpenStudy (magbak):

It is C

OpenStudy (magbak):

Point slope is\[y-y_1= m(x-x_1)\]

OpenStudy (magbak):

Got it??

OpenStudy (mikezack123):

Got it! thanks!

OpenStudy (magbak):

You are welcom need any more help ??

OpenStudy (mikezack123):

Thats all for right now, thanks

OpenStudy (magbak):

Ok if you need any thing just message me ok or tag me. LOL :)

OpenStudy (mikezack123):

k :)

OpenStudy (mikezack123):

@magbak i got my scores back, i got the second one wrong can you help me again?

OpenStudy (magbak):

Yes but is that one that I did or not? and what corse are you taking?

OpenStudy (mikezack123):

algebra 1, you did the second question, and nevermind i did it myself, using yours, and @phi help, thanks :)

OpenStudy (magbak):

So I knew my answer was right.

OpenStudy (magbak):

Can you reset it because I am taking algebra 2 and I can reset assignments

OpenStudy (mikezack123):

no the the second one that YOU did was wrong phi did the 3rd

OpenStudy (magbak):

Oh I sorry. But can you reset it or not